# The satisfiability threshold and solution space of random uniquely extendable constraint satisfaction problems

### v1

@@ 1 Introduction
Let $k\geq 1$ be an integer. A $k$-ary constraint $\psi$ is called uniquely extendable, if for every $1\leq i\leq k$ and for every $\sigma_{1},\ldots,\sigma_{i-1},\sigma_{i+1},\ldots,\sigma_{k}\in\Omega^{k-1}$, there is a unique $\tau\in\Omega$ such that $\psi(\sigma_{1},\ldots,\sigma_{i-1},\tau,\sigma_{i+1},\ldots,\sigma_{k})=1$.
@@ 1.1 Main results
A $k$-ary constraint $\psi$ on $\Omega$ is commutative if, for all $\sigma\in\Omega^{k}$ and $\pi\in\Pi([k])$,
@@
A $k$-ary constraint $\psi$ is reducible if, for all $d\in\{0,\ldots,k\}$, $\sigma,\sigma^{\prime}\in\Omega^{d}$, and $\eta,\eta^{\prime}\in\Omega^{k-d}$ such that
@@
For any $k$-ary UE constraint function $\psi$, let $f_{\psi}:\Omega^{k-1}\to\Omega$ be the unique function that $\sigma f_{\psi}(\sigma)\models\psi$ for all $\sigma\in\Omega^{k-1}$.
@@
    Let $F_{0}(\Psi)$ be the set of the identity function on $\Omega$. That is, $F_{0}(\Psi)$ contains a single function which maps $\sigma$ to $\sigma$ for every $\sigma\in\Omega$.
@@
Let $\Psi\subseteq\Lambda_{k,r}$. We say that $\Psi$ is reducible if, for all $d\in\{0,\ldots,k\}$, $\sigma,\sigma^{\prime}\in\Omega^{d}$, $\eta,\eta^{\prime}\in\Omega^{k-d}$, and $\psi,\psi^{\prime}\in\Psi$ such that
@@
    If $m>(d_{k}/k+\varepsilon)n$ and $Q_{\pi}\neq\emptyset$, then a.a.s. the solutions of $H_{n}(\pi,k,m)$ are partitioned into $h=|Q_{\pi}|$ clusters that satisfy properties (ii)–(iv) in part (b).
@@
Note that if $Q_{\pi}\neq\emptyset$ then $\pi$ is supported on a single $\psi$ by the reducibility of $\operatorname{supp}(\pi)$ if $k\geq 4$ or the symmetry of $F_{2}(\operatorname{supp}(\pi))$ if $k=3$, which is assumed by the hypothesis of the theorem.
We expect that the phase transition of RSB occurs when $m/n$ is around $d_{k}/k$ when $\operatorname{supp}(\pi)$ satisfies the same conditions as in Theorem 1.13. Indeed, it follows easily from our results in this paper (see Remark 1.22 below) when $r$ is the product of distinct prime numbers. However, for general $r$ we did not manage to prove that the replica symmetry holds until reaching the satisfiability threshold, and thus we leave it as a conjecture; see Conjecture 1.21.
@@ 1.2 An overview of the proofs
Suppose $k\geq 3$ and suppose $\Psi\subseteq\Lambda_{k,r}$ is such that every $\psi\in\Psi$ is commutative. Then the following are equivalent
@@
Suppose $k\geq 4$ and suppose $\Psi\subseteq\Lambda_{k,r}$ is such that every $\psi\in\Psi$ is commutative. Then the following are equivalent
@@
    The assumption $k\geq 4$ in Theorem 1.17 is necessary. In fact, it is easy to show that every $3$-ary UE commutative constraint function is reducible (we leave this as an easy exercise to the reader). However, not all these constraint functions are equivalent to group constraint functions. A concrete example for $(k,q)=(3,4)$ was given by Connamacher and Molloy where $\Psi$ contains three constraint functions, each as a group constraint, but $F_{2}(\Psi)$ is asymmetric. We provide a family of $3$-ary UE commutative constraint functions that are not equivalent to any group constraint function in section 5. The same construction indeed also provides a family of $k$-ary non-reducible UE constraint functions for $k\geq 4$.
@@
    The special case of $\Psi=\{\psi\}$ of Theorems 1.16 and 1.17 can be proved using known results from quasigroup theory, and we sketch these arguments in the Appendix. It is possible that the quasigroup results we rely on can be further generalised to handle the full versions of Theorems 1.16 and 1.17. We have not pursued this direction. Instead, we provide self-contained proofs of these theorems without using quasigroup theory; these proofs may be of independent interest.
@@ 1.3 More conjectures and future work
Finally, we make the following conjecture on the phase transition of replica symmetry breaking at $d_{k}$. It has been shown that the “replica symmetry breaking” (or “condensation”) occurs at density $d_{k}$ for random linear equations over finite fields. We conjecture that the same holds for $H_{n}(\pi,k,m)$, at least when $\operatorname{supp}(\pi)$ is reducible.
@@
    ${\mathbb{P}}[x_{u}=\sigma,x_{v}=\tau]=1/q^{2}+o(1),\quad\text{for all $\sigma,\tau\in\Omega$}.$ (2)
@@
    If $m>(d_{k}/k+\varepsilon)n$ and $Q_{\psi}\neq\emptyset$ where $Q_{\psi}$ is the set of constant solutions of $\psi$, then, for a uniform random solution $x$ of $H_{n}(\pi,k,m)$, property (2) does not hold a.a.s.
@@
Part (b) of the conjecture follows easily from Our Theorem 1.13 (c). By our Theorems 1.161.17, and the RS property of the solution space of random linear equations over finite fields [ayre2020satisfiability, eq. (2.5)] if $r$ is a product of distinct prime numbers. However, this approach of reduction does not work if for instance $r=p^{k}$ for some prime $p$ and $k\geq 2$, since the solutions of linear equations over $GF(r)$ is not bijectively mapped to the solutions over ${\mathbb{Z}}_{r}$.
@@
By Lemma 2.1, the rows of $B^{\prime}$ are linearly independent a.a.s. over $\mathbb{F}_{p}$. By the definiton of the 2-core of $B^{\prime}$, each row in $R_{n}$ is supported on $C_{n}$. Consequently, we may assume the event that the rows of $B^{\prime}_{\text{2-core}}$ are linearly independent over $\mathbb{F}_{p}$, which holds a.a.s. Then there is an $|R_{n}|\times|R_{n}|$ submatrix $D$ of $B^{\prime}_{\text{2-core}}$ that is invertible over $\mathbb{F}_{p}$. Similarly as in the subcritical case, we can conclude that $D$ is invertible over $\mathbb{Z}_{q}$. Therefore the number of solutions to $Dx=b|_{R_{n}}$ over $\mathbb{Z}_{q}$ is $|\mathbb{Z}_{q}|^{|C_{n}|-|R_{n}|}=q^{O(\varepsilon_{n}n)}$ by Lemma 2.1. The claimed assertion now follows since every solution to $B^{\prime}_{\text{2-core}}\,x=b|_{R_{n}}$ must be a solution to $Dx=b|_{R_{n}}$. ∎
@@ 3.1 Proof of Parts (a,b)
Suppose $k\geq 3$ and $m=dn$ for some $d>d_{k}^{\star}$. Then the expected number of vertices in flippable cycles of $B$ is $O(1)$.
First consider the case where $Bx=b$ is a system of equation over a cyclic group $\mathbb{Z}_{q}$, where $q$ is a power of some prime $p$. By definition of the solution clusters and Lemma 3.3, the number of clusters $N$ is the number of solutions to $B_{2-\text{core}}x=b$ up to a factor of $q^{\log n}$, as two such solutions might be contained in the same cluster, if they differ only on the set of variables that are contained in flippable cycles of $B_{\text{2-core}}$ — by Lemma 3.3 there are a.a.s. at most $\log n$ such variables. Let $m^{*}\times n^{*}$ be the dimensions of $B_{\text{2-core}}$. If we consider $B_{2-\text{core}}$ as a binary matrix over $\mathbb{F}_{p}$, Lemma 2.1 shows that the rows of $B_{\text{2-core}}$ are linearly independent. There is therefore an $m^{*}\times m^{*}$ binary invertible submatrix $D$ of $B_{\text{2-core}}$. Since $\det(D)\neq 0\pmod{p}$, we have $\det(D)\neq 0\pmod{q}$ and thus $D$ is invertible over $\mathbb{Z}_{q}$. Therefore the number of $2$-core solutions is $q^{n^{*}-m^{*}}$, and thus $N=q^{n^{*}-m^{*}+O(\log n)}$, where $N$ is the number of clusters. By [ayre2020satisfiability, Theorem 1.6] for $n^{*}/n$ and $m^{*}/m$, we get
@@ 3.2 Proof of part c
We prove by induction on $i$. Suppose $i=1$, and suppose $d>0$ is such that $\pi_{k}(d_{k})\leq\pi_{k}(d)<\pi_{k}(d_{k})+\delta/2$. Let $\varepsilon_{n}$ be chosen to satisfy Lemma 2.1, and let $m^{\prime}=(d_{k}/k-\varepsilon_{n})n$ and $m^{\prime\prime}=m-m^{\prime}$. Let $B^{\prime}$ be the submatrix consisting of the first $m^{\prime}$ rows of $B$, and let $B^{\prime\prime}$ be the submatrix consisting of the last $m^{\prime\prime}$ rows of $B$. Let $R_{n}$ and $C_{n}$ be the rows and columns of $B^{\prime}_{\text{2-core}}$ respectively, and let $R^{\prime}_{n}$ be the set of rows of $B^{\prime\prime}$ whose entries are zero outside of $C_{n}$. Let $\Sigma$ be the set of solutions to $B_{\text{2-core}}\,x=b\mid_{R_{n}}$. These are exactly the same definitions as in the proof of Theorem 1.15. Thus, we immediately have that a.a.s. $|R^{\prime}_{n}|=\Omega(m^{\prime\prime})$ by (3), and $|\Sigma|=q^{O(\varepsilon_{n}n)}$ by Claim 2.2.
@@
Let $\hat{x}$ satisfy $B\hat{x}=b$ and suppose for a contradiction that $\hat{x}$ is not cycle equivalent to any solution in $Q_{\Psi}$ with respect to the 2-core of $B$. Then, by (P2), $\hat{x}$ differs from each $\sigma^{n}$, where $\sigma\in Q_{\pi}$, on at least $\delta n$ variables in the 2-core of $B$. Moreover, $n_{*}(B)/n-n_{*}(B^{\prime})/n\leq 3\delta n/4$ by (P3), and the variables of the 2-core of $B^{\prime}$ is a subset of the variables in the 2-core of $B$. It follows then that $\hat{x}$ must differ from each $\sigma^{n}$, where $\sigma\in Q_{\pi}$, on at least $\delta n/4$ variables of the 2-core of $B^{\prime}$ a.a.s. By (P1), the solution $\hat{x}$ is not cycle equivalent to a constant solution on the 2-core of $B^{\prime}$. This is a contradiction since we have shown that a.a.s. all solutions of $Bx=b$ are cycle equivalent, with respect to the 2-core of $B^{\prime}$, to a constant solution. Hence, $\hat{x}$ is cycle equivalent, with respect to $B_{\text{2-core}}$, to $\sigma^{n}$ for some $\sigma\in Q_{\pi}$.
@@ 4 Characterization of group constraint functions for $k\geq 3$
For the converse direction, suppose every function in $F_{2}(\Psi)$ is symmetric. Let $\psi_{1},\psi_{2}\in\Psi$. Let $d\in[k-1]$ and let $\sigma_{1},\sigma_{2}\in\Omega^{d}$, $\eta_{1},\eta_{2}\in\Omega^{k-d}$ be such that $\sigma_{1}\eta_{1},\sigma_{2}\eta_{1}\models\psi_{1}$ and $\sigma_{1}\eta_{2}\models\psi_{2}$. Let $\eta_{1}^{\prime},\eta_{2}^{\prime}\in\Omega^{k-d-1}$ be the prefixes of $\eta_{1},\eta_{2}$ respectively. Since $\sigma_{1}\eta_{1},\sigma_{2}\eta_{1}\models\psi_{1}$, we have $\sigma_{1}\eta_{1}^{\prime}f_{\psi_{1}}(\sigma_{1}\eta_{1}^{\prime})=\sigma_{1}\eta_{1}$ and $\sigma_{2}\eta_{1}^{\prime}f_{\psi_{1}}(\sigma_{2}\eta_{1}^{\prime})=\sigma_{2}\eta_{1}$, and in particular $f_{\psi_{1}}(\sigma_{1}\eta_{1}^{\prime})=f_{\psi_{1}}(\sigma_{2}\eta_{1}^{\prime})$. Now, since every function in $F_{2}(\Psi)$ is symmetric, for any $\eta_{3},\ldots\eta_{k-1}\in\Omega^{k-1}$,
@@
Thus, by unique extendability of $\psi_{1}$, we get $f_{\psi_{2}}(\sigma_{2}\eta_{2}^{\prime})=f_{\psi_{2}}(\sigma_{1}\eta_{2}^{\prime})$. Since $\sigma_{1}\eta_{2}\models\psi_{2}$, we have $\sigma_{1}\eta_{2}^{\prime}f_{\psi_{2}}(\sigma_{1}\eta_{2}^{\prime})=\sigma_{1}\eta_{2}$, and so $\eta_{2}=\eta_{2}^{\prime}f_{\psi_{2}}(\sigma_{1}\eta_{2}^{\prime})=\eta_{2}^{\prime}f_{\psi_{2}}(\sigma_{2}\eta_{2}^{\prime})$. Therefore $\sigma_{2}\eta_{2}=\sigma_{2}\eta_{2}^{\prime}f_{\psi_{2}}(\sigma_{2}\eta_{2}^{\prime})\models\psi_{2}$. ∎

### v2

@@ 1 Introduction
Let $k\geq 1$ be an integer. A $k$-ary constraint function $\psi$ is called uniquely extendable, if for every $1\leq i\leq k$ and for every $\sigma_{1},\ldots,\sigma_{i-1},\sigma_{i+1},\ldots,\sigma_{k}\in\Omega$, there is a unique $\tau\in\Omega$ such that $\psi(\sigma_{1},\ldots,\sigma_{i-1},\tau,\sigma_{i+1},\ldots,\sigma_{k})=1$.
@@ 1.1 Main results
A $k$-ary constraint function $\psi$ on $\Omega$ is commutative if, for all $\sigma\in\Omega^{k}$ and $\pi\in\Pi([k])$,
@@
A $k$-ary constraint function $\psi$ is reducible if, for all $d\in\{0,\ldots,k-1\}$, $\sigma,\sigma^{\prime}\in\Omega^{d}$, and $\eta,\eta^{\prime}\in\Omega^{k-d}$ such that
@@
For any $k$-ary UE constraint function $\psi$, let $f_{\psi}:\Omega^{k-1}\to\Omega$ be the unique function such that $\sigma f_{\psi}(\sigma)\models\psi$ for all $\sigma\in\Omega^{k-1}$.
@@
    Let $F_{0}(\Psi)$ be the singleton set containing the identity function on $\Omega$. That is, $F_{0}(\Psi)$ contains a single function which maps $\sigma$ to $\sigma$ for every $\sigma\in\Omega$.
@@
Let $\Psi\subseteq\Lambda_{k,r}$. We say that $\Psi$ is reducible if, for all $d\in\{0,\ldots,k-1\}$, $\sigma,\sigma^{\prime}\in\Omega^{d}$, $\eta,\eta^{\prime}\in\Omega^{k-d}$, and $\psi,\psi^{\prime}\in\Psi$ such that
@@
    If $m>(d_{k}/k+\varepsilon)n$ and $Q_{\pi}\neq\emptyset$, then a.a.s. the solutions of $H_{n}(\pi,k,m)$ are partitioned into $|Q_{\pi}|$ clusters that satisfy properties (ii)–(iv) in part (b).
@@
Note that if $Q_{\pi}\neq\emptyset$ then $\pi$ is supported on a single $\psi$ by the reducibility of $\operatorname{supp}(\pi)$ if $k\geq 4$ or the symmetry of $F_{2}(\operatorname{supp}(\pi))$ if $k=3$, which is assumed by the hypothesis of the theorem. We leave this observation as a simple exercise. Our proof of Theorem 1.13 does not rely on this fact.
We expect that the phase transition of RSB occurs when $m/n$ is around $d_{k}/k$ if $\operatorname{supp}(\pi)$ satisfies the same conditions as in Theorem 1.13. Indeed, it follows easily from our results in this paper (see Remark 1.22 below) when $r$ is the product of distinct prime numbers. However, for general $r$ we did not manage to prove that the replica symmetry holds until reaching the satisfiability threshold, and thus we leave it as a conjecture; see Conjecture 1.21.
@@ 1.2 An overview of the proofs
Suppose $k\geq 3$ and suppose $\Psi\subseteq\Lambda_{k,r}$ is such that every $\psi\in\Psi$ is commutative. Then the following are equivalent.
@@
Suppose $k\geq 4$ and suppose $\Psi\subseteq\Lambda_{k,r}$ is such that every $\psi\in\Psi$ is commutative. Then the following are equivalent.
@@
    The assumption $k\geq 4$ in Theorem 1.17 is necessary. In fact, it is easy to show that every $3$-ary UE commutative constraint function is reducible (we leave this as an easy exercise to the reader). However, not all these constraint functions are equivalent to group constraint functions. A concrete example for $(k,q)=(3,4)$ was given by Connamacher and Molloy where $\Psi$ contains three constraint functions, each as a group constraint, but $F_{2}(\Psi)$ is not symmetric. We provide a family of $3$-ary UE commutative constraint functions that are not equivalent to any group constraint function in section 5. The same construction indeed also provides a family of $k$-ary non-reducible UE constraint functions for $k\geq 4$.
@@
    The special case of $\Psi=\{\psi\}$ of Theorems 1.16 and 1.17 can be proved using known results from quasigroup theory, and we sketch these arguments in the Appendix. It is possible that the quasigroup results we rely on can be further generalised to handle the full versions of Theorems 1.16 and 1.17. We have not pursued this direction. Instead, we provide self-contained proofs of these theorems without using quasigroup theory, and we believe that these proofs are of independent interest.
@@ 1.3 More conjectures and future work
Note. The setting of this conjecture permits a random system of group equations of form (1), where different equations can be over different groups, provided that all the groups have the same order. This flexibility is motivated by the fact that the SAT threshold in Theorem 1.15 is independent of the group structure of $G$. However, our current proof crucially relies on the assumption that all equations are defined over a single common group.
Finally, we make the following conjecture on the phase transition of replica symmetry breaking at $d_{k}$. It has been shown that the “replica symmetry breaking” (or “condensation”) occurs at density $d_{k}$ for random linear equations over finite fields. We conjecture that the same holds for $H_{n}(\pi,k,m)$, at least when $\operatorname{supp}(\pi)$ is reducible. Recall the definition of $Q_{\pi}$ before Theorem 1.13.
@@
    ${\mathbb{P}}[x_{u}=\sigma,x_{v}=\tau]=1/r^{2}+o(1),\quad\text{for all $\sigma,\tau\in\Omega$}.$ (2)
@@
    If $m>(d_{k}/k+\varepsilon)n$ and $Q_{\pi}\neq\emptyset$, then, for a uniform random solution $x$ of $H_{n}(\pi,k,m)$, property (2) does not hold a.a.s.
@@
Part (b) of the conjecture follows easily from Our Theorem 1.13 (c). By our Theorems 1.161.17, and the RS property of the solution space of random linear equations over finite fields [ayre2020satisfiability, eq. (2.5)], part (a) of the conjecture follows immediately if $r$ is a product of distinct prime numbers. However, this approach of reduction does not work if for instance $r=p^{k}$ for some prime $p$ and $k\geq 2$, since the solutions of linear equations over $GF(r)$ is not bijectively mapped to the solutions over ${\mathbb{Z}}_{r}$.
@@
By Lemma 2.1, the rows of $B^{\prime}$ are linearly independent a.a.s. over $\mathbb{F}_{p}$. By the definiton of the 2-core of $B^{\prime}$, each row in $R_{n}$ is supported on $C_{n}$. Consequently, we may assume the event that the rows of $B^{\prime}_{\text{2-core}}$ are linearly independent over $\mathbb{F}_{p}$, which holds a.a.s. Then there is an $|R_{n}|\times|R_{n}|$ submatrix $D$ of $B^{\prime}_{\text{2-core}}$ that is invertible over $\mathbb{F}_{p}$. Similarly as in the subcritical case, we can conclude that $D$ is invertible over $\mathbb{Z}_{q}$. Therefore the number of solutions to $B^{\prime}_{\text{2-core}}\,x=b|_{R_{n}}$ is $|\mathbb{Z}_{q}|^{|C_{n}|-|R_{n}|}$, which is $q^{O(\varepsilon_{n}n)}$ by Lemma 2.1. ∎
@@ 3.1 Proof of Parts (a,b)
Suppose $k\geq 3$ and $m=dn/k$ for some $d>d_{k}^{\star}$. Then the expected number of vertices in flippable cycles of $B$ is $O(1)$.
First consider the case where $Bx=b$ is a system of equations over a cyclic group $\mathbb{Z}_{q}$, where $q$ is a power of some prime $p$. By definition of the solution clusters and Lemma 3.3, the number of clusters $N$ is the number of solutions to $B_{2-\text{core}}x=b$ up to a factor of $q^{\log n}$, as two such solutions might be contained in the same cluster, if they differ only on the set of variables that are contained in flippable cycles of $B_{\text{2-core}}$ — by Lemma 3.3 there are a.a.s. at most $\log n$ such variables. Let $m^{*}\times n^{*}$ be the dimensions of $B_{\text{2-core}}$. If we consider $B_{2-\text{core}}$ as a binary matrix over $\mathbb{F}_{p}$, Lemma 2.1 shows that the rows of $B_{\text{2-core}}$ are linearly independent. There is therefore an $m^{*}\times m^{*}$ binary invertible submatrix $D$ of $B_{\text{2-core}}$. Since $\det(D)\neq 0\pmod{p}$, we have $\det(D)\neq 0\pmod{q}$ and thus $D$ is invertible over $\mathbb{Z}_{q}$. Therefore the number of $2$-core solutions is $q^{n^{*}-m^{*}}$, and thus $N=q^{n^{*}-m^{*}+O(\log n)}$, where $N$ is the number of clusters. By [ayre2020satisfiability, Theorem 1.6] for $n^{*}/n$ and $m^{*}/m$, we get
@@ 3.2 Proof of part c
We prove by induction on $i$. Suppose $i=1$, and suppose $d>0$ is such that $\pi_{k}(d_{k})\leq\pi_{k}(d)<\pi_{k}(d_{k})+\delta/2$. Let $\varepsilon_{n}$ be chosen to satisfy Lemma 2.1, and let $m^{\prime}=(d_{k}/k-\varepsilon_{n})n$ and $m^{\prime\prime}=m-m^{\prime}$. Let $B^{\prime}$ be the submatrix consisting of the first $m^{\prime}$ rows of $B$, and let $B^{\prime\prime}$ be the submatrix consisting of the last $m^{\prime\prime}$ rows of $B$. Let $R_{n}$ and $C_{n}$ be the rows and columns of $B^{\prime}_{\text{2-core}}$ respectively, and let $R^{\prime}_{n}$ be the set of rows of $B^{\prime\prime}$ whose entries are zero outside of $C_{n}$. Let $\Sigma$ be the set of solutions to $B_{\text{2-core}}^{\prime}\,x=b\mid_{R_{n}}$. These are exactly the same definitions as in the proof of Theorem 1.15. Thus, we immediately have that a.a.s. $|R^{\prime}_{n}|=\Omega(m^{\prime\prime})$ by (3), and $|\Sigma|=q^{O(\varepsilon_{n}n)}$ by Claim 2.2.
@@
Let $\hat{x}$ satisfy $B\hat{x}=b$ and suppose for a contradiction that $\hat{x}$ is not cycle equivalent to any $\sigma^{n}$, with respect to the 2-core of $B$, where $\sigma\in Q_{\pi}$. Then, by (P2), $\hat{x}$ differs from each $\sigma^{n}$, where $\sigma\in Q_{\pi}$, on at least $\delta n$ variables in the 2-core of $B$. Moreover, $n_{*}(B)/n-n_{*}(B^{\prime})/n\leq 3\delta n/4$ by (P3), and the variables of the 2-core of $B^{\prime}$ is a subset of the variables in the 2-core of $B$. It follows then that $\hat{x}$ must differ from each $\sigma^{n}$, where $\sigma\in Q_{\pi}$, on at least $\delta n/4$ variables of the 2-core of $B^{\prime}$ a.a.s. By (P1), the solution $\hat{x}$ is not cycle equivalent to a constant solution on the 2-core of $B^{\prime}$. This is a contradiction since we have shown that a.a.s. all solutions of $Bx=b$ are cycle equivalent, with respect to the 2-core of $B^{\prime}$, to a constant solution. Hence, $\hat{x}$ is cycle equivalent, with respect to $B_{\text{2-core}}$, to $\sigma^{n}$ for some $\sigma\in Q_{\pi}$.
@@ 4 Characterization of group constraint functions for $k\geq 3$
For the converse direction, suppose every function in $F_{2}(\Psi)$ is symmetric. Let $\psi_{1},\psi_{2}\in\Psi$. Let $d\in\{0,\ldots,k-1\}$ and let $\sigma_{1},\sigma_{2}\in\Omega^{d}$, $\eta_{1},\eta_{2}\in\Omega^{k-d}$ be such that $\sigma_{1}\eta_{1},\sigma_{2}\eta_{1}\models\psi_{1}$ and $\sigma_{1}\eta_{2}\models\psi_{2}$. Let $\eta_{1}^{\prime},\eta_{2}^{\prime}\in\Omega^{k-d-1}$ be the prefixes of $\eta_{1},\eta_{2}$ respectively. Since $\sigma_{1}\eta_{1},\sigma_{2}\eta_{1}\models\psi_{1}$, we have $\sigma_{1}\eta_{1}^{\prime}f_{\psi_{1}}(\sigma_{1}\eta_{1}^{\prime})=\sigma_{1}\eta_{1}$ and $\sigma_{2}\eta_{1}^{\prime}f_{\psi_{1}}(\sigma_{2}\eta_{1}^{\prime})=\sigma_{2}\eta_{1}$, and in particular $f_{\psi_{1}}(\sigma_{1}\eta_{1}^{\prime})=f_{\psi_{1}}(\sigma_{2}\eta_{1}^{\prime})$. Now, since every function in $F_{2}(\Psi)$ is symmetric, for any $\eta_{3},\ldots\eta_{k-1}\in\Omega^{k-1}$,
@@
Thus, by unique extendability of $\psi_{1}$, we get $f_{\psi_{2}}(\sigma_{2}\eta_{2}^{\prime})=f_{\psi_{2}}(\sigma_{1}\eta_{2}^{\prime})$. Since $\sigma_{1}\eta_{2}\models\psi_{2}$, we also have $\sigma_{1}\eta_{2}^{\prime}f_{\psi_{2}}(\sigma_{1}\eta_{2}^{\prime})=\sigma_{1}\eta_{2}$, and so $\eta_{2}=\eta_{2}^{\prime}f_{\psi_{2}}(\sigma_{1}\eta_{2}^{\prime})=\eta_{2}^{\prime}f_{\psi_{2}}(\sigma_{2}\eta_{2}^{\prime})$. Therefore $\sigma_{2}\eta_{2}=\sigma_{2}\eta_{2}^{\prime}f_{\psi_{2}}(\sigma_{2}\eta_{2}^{\prime})\models\psi_{2}$. ∎